OpenStudy (anonymous):

Given a cube with a volume of 27 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? 3cm3 6cm3 9cm3 12cm3

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (mathstudent55):

Find the formulas for the volumes of a cube and a pyramid and write them here. Then we'll continue.

OpenStudy (anonymous):

idk this stuff

OpenStudy (anonymous):

Can you just help me through it?

OpenStudy (mathstudent55):

You must be able to find the formulas you need. Do you have a text book?

OpenStudy (mathstudent55):

Ok, here are the formulas. \(\Large V_{cube} = s^3\) where s = side of cube \(\Large V_{pyramid} = \dfrac{1}{3}Bh\) where B = area of the base,. and h = height of the pyramid

OpenStudy (anonymous):

No I dont have a text book

OpenStudy (mathstudent55):

Here is a cube with side s. The volume of the cube is \(s^3\).